OpenStudy (anonymous):

Could someone please help me with the quadratic formula for 4x2 = - 9x - 4? I'm not sure of my process in the end.

OpenStudy (mathstudent55):

First, we need the quadratic formula. \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (mathstudent55):

Now we need to change your equation into the correct form to use it in the quadratic formula.

OpenStudy (mathstudent55):

\(4x^2 = -9x - 4\) We need everything on the left side, so we add to both sides 9x and 4. \(4x^2 + 9x + 4 = 0\) Do you follow this so far?

OpenStudy (anonymous):

Yes, I do. Thanks! but I know that later on in the problem there has to be a negative 2 (because of the options given) and I'm not sure where that came from...